OpenStudy (smurfy14):

y=cos(-0.4x) has a horizontal stretch of..? PLEASE HELP

OpenStudy (anonymous):

you mean period?

OpenStudy (smurfy14):

oh yes

OpenStudy (anonymous):

are you looking for the period of \[\cos(-.4x)\]?

OpenStudy (smurfy14):

yes

OpenStudy (smurfy14):

im looking for something like "it stretches X..."

OpenStudy (anonymous):

period of \[\cos(x)\] is \[2\pi\] so period of \[\cos(bx)\] is \[\frac{2\pi}{b}\]

OpenStudy (anonymous):

in your case you get \[\frac{2\pi}{.4}=5\pi\]

OpenStudy (smurfy14):

no heres an example: y=-3cos1/2x has a horizontal stretch of 2 but i cant figure out what it would be for this problem

OpenStudy (anonymous):

so period goes from 2 pi to 5 pi

OpenStudy (anonymous):

oooooooooooooooh silly way to word the problem. in your case \[\cos(\frac{1}{2}x)\] teh "stretch" was 2

OpenStudy (anonymous):

in other words the reciprocal of coefficient. the reciprocal of .4 is 5, so stretch is 5

OpenStudy (smurfy14):

ohhhkay so you dont have to include the whole thing like 2/5=5/2 you just take the numerator?